OpenStudy (anonymous):

Anna builds a rectangular prism storage container with dimensions of 8 inches by 6 inches by 12 inches. She builds a second container with a greater volume by increasing the dimensions of the rectangular prism to 16 inches by 12 inches by 24 inches. How will the volume of the second container she builds compare to the volume of the first one?

OpenStudy (anonymous):

If you double the size of a rectangle, the larger rectangle is 2^2 the size of the smaller. If you double the size of a box. the larger box is 2^3 the volume of the smaller. Your smaller box volume is V1 = 8*6*12 Your larger box is v2 = 16*12*24 The ratio of V2 to V1 is V2/V1

OpenStudy (anonymous):

stupid question, what does ^ mean

OpenStudy (anonymous):

"to the power of". But ^2 and ^3 have special names. Instead of 2^2 being called "2 to the power of 2" (even though it is), it's simply called "2 squared". Likewise, "^3" is "cubed", so that 2^3 is "2 cubed".

OpenStudy (anonymous):

wow.. thanks a ton :P

OpenStudy (anonymous):

No problem! Hop I've helped! :-)